The overall equation for one method of manufacturing
hydrogen from methane is
CH4 + 2H2O → 4H2 + CO2
In one batch, 5.12 tonnes of hydrogen were obtained from
13.6 tonnes of methane.
What is the percentage yield of hydrogen for this batch?
A 37.6
B 42.5
C 54.4
D 75.3
1
Expert's answer
2020-10-08T13:57:28-0400

m1(H2) = 5.12 tones

m(CH4) = 13.6 tonnes = 13 600 000 g

n(CH4) = m(CH4)/M(CH4) = 13 600 000/(12+1*4)= 850 000 mole


850 000 mole x mole

CH4 + 2H2O → 4H2 + CO2

1 mole 4 mole


850 000/1 = x/4


x = (850 000*4)/1 = 3 400 000 mole = n(H2)

m2(H2) = n(H2)*M(H2) = 3 400 000 * (1+1) = 6 800 000 g = 6.8 tones


percentage yield of hydrogen = m1(H2)/m2(H2) * 100% = 5.12/6.8 * 100% = 75.3%


answer D 75.3